OpenStudy (anonymous):

Suppose y varies inversely with x. Find an equation that relates x and y given the following. x=8 when y=5

OpenStudy (anonymous):

y varies inversely with x this means: y = (1/x)k where k is a constant 5 = k/8 k = 8(5) = 40 y = 40/x